# Stochastic LQR Design With Disturbance Preview

### v3

@@ I Introduction
There is a large literature on finite preview control for stochastic LQR/LQG and its closely related $H_{2}$ formulation. These results exist in both continuous-time [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and discrete-time [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. A common approach in discrete-time is to construct an augmented system with a chain of delays to store the disturbance preview information. This technique simplifies the preview control design as the standard LQR/LQG/$H_{2}$ solutions can be applied to this augmented model [10, 19, 20, 12, 13, 22, 23, 24, 25, 15, 26, 16, 17, 27, 28]. The work in [14] is an exception as it use calculus of variations to directly derive necessary conditions for optimality.
@@ II Notation
Let $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times m}$ denote the sets of real $n\times 1$ vectors and $n\times m$ matrices, respectively. The superscript $\top$ denote the matrix transpose. Moreover, if $M\in\mathbb{R}^{n\times n}$ then $M\succ 0$ denotes that $M$ is a symmetric positive definite matrix. Similarly, we use $\succeq$, $\prec$, and $\preceq$ for positive semidefinite, negative definite, and negative semidefinite matrices, respectively. A matrix $A\in\mathbb{R}^{n\times n}$ is said to be Schur if the spectral radius is $<1$. We also introduce random variables in the problem formulation and use $\mathbb{E}$ to denote the expectation. Finally, we denote the Kronecker delta $\delta_{ij}$ where $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise.
@@ III Problem Formulation
$i_{t}$ :-$\coloneq\{x_{0},\dots,x_{t},w_{0},\dots,w_{t+p}\}.:$ (7)
@@
Consider the LTI plant (6) with $A$ nonsingular, $(A,B)$ stabilizable and cost function in (56) with $Q\succeq 0$, $R\succ 0$ and $(A,Q)$ detectable. Find a sequence of policies $K\coloneq\{K_{0},K_{1}\dots\}$ to stabilize the plant and solve:
@@ IV Principle of Optimality
$W_{T}(i_{T})$ :-$\coloneq\mathbb{E}[g_{T}(\boldsymbol{x}_{T})\,|\,i_{T}],:$
$W_{t}(i_{t})$ :-$\coloneq\min_{u}\mathbb{E}\left[g_{t}(x_{t},u_{t})+W_{t+1}(i_{t+1})\,\middle|\,i_{t},u_{t}=u\right]:$
[…]
@@
$V_{t}^{K_{t:T-1}}(i_{t})\geq\mathbb{E}\left[\vphantom{\Big{|}}\mathbb{E}\left[g_{t}(x_{t},u_{t})+W_{t+1}(i_{t+1})\,\middle|\,i_{t},u_{t}=u\right]\,\middle|\,i_{t}\right].$ (17)
@@
The policy that achieves $J^{*}(i_{0})$ from will be deterministic if we use this choice for $t=0$ to $T-1$. ∎
@@ V-A FH Stochastic LQR with Preview
$K_{x,t}$ :-$\coloneq H_{t}^{-1}B_{u,t}^{\top}P_{t+1}A_{t},:$ (22)
$K_{w,t}$ :-$\coloneq H_{t}^{-1}B_{u,t}^{\top}P_{t+1}B_{w,t},:$ (23)
$K_{v,t}$ :-$\coloneq H_{t}^{-1}B_{u,t}^{\top}.:$ (24)
@@ V-B IH Stochastic LQR with Preview
$K_{x}$ :-$\coloneq H^{-1}B_{u}^{\top}PA,:$ (37)
$K_{w}$ :-$\coloneq H^{-1}B_{u}^{\top}PB_{w},:$ (38)
$K_{v}$ :-$\coloneq H^{-1}B_{u}^{\top}.:$ (39)
@@
Next, use the definition of $u_{t}^{*}$ in (40) to replace $K_{x}x_{t}$ with $-u_{t}^{*}-K_{w}w_{t}-K_{v}v_{t+1}$. The per-step cost is thus given by
@@
Here we have used $K_{w}=K_{v}PB_{w}$ to simplify the expression further. Note that the optimal IH cost does not depend on the initial condition, i.e. $J_{\infty,p}^{*}(i_{0})=J_{\infty,p}^{*}$. The first term in the optimal cost, trace$[PB_{w}B_{w}^{\top}]$, is the cost that would be achieved using only state feedback. The second term gives the cost reduction obtained by using the current disturbance measurement and $p$ steps of preview.
@@ V-C Relation to Optimal Noncausal Controller
$i_{t}$ :-$\coloneq\{x_{0},\dots,x_{t},w_{0},w_{1},\ldots\}.:$ (48)
@@
Consider the LTI plant (6) with $A$ nonsingular, $(A,B)$ stabilizable and cost function in (49) with $Q\succeq 0$, $R\succ 0$ and $(A,Q)$ detectable. Also assume $w\in\ell_{2}$. Find a sequence of deterministic policies $K\coloneq\{k_{0},k_{1}\dots\}$ to stabilize the plant and solve:
@@
This result, as stated, is a special case of Theorem 1 in [34]. It was stated in [34] for a more general LQR cost including a cross term $2x_{t}^{\top}Sw_{t}$. [^4] The proof given in [34] was for signals defined for $t$ from $\infty$- ∞ to +$\infty$+ ∞. However, the optimal noncausal controller is the same for signals defined from $t=0$ to $\infty$. Note that the full preview design (Problem 3) was stated with a deterministic formulation while finite preview design (Problem 2) was derived with stochastic formulation. However, we show next that the controller with finite preview $p<\infty$ converges, in the limit as $p\to\infty$, to the optimal noncausal controller for any fixed disturbance $w\in\ell_{2}$.
@@
Next, we show that $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\to 0$ uniformly in $t$ as $p\to\infty$. The difference between the two commands is:
$u_{t}^{nc}-u_{t}^{p}=-K_{v}\cdot\sum_{j=t+p+1}^{\infty}\left[\hat{A}^{\top}\right]^{j-t}PB_{w}w_{j}$ (55)
We can bound this difference as $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\leq a_{p}\cdot b$ where
$a_{p}:=\left\|\sum_{l=p+1}^{\infty}[\hat{A}^{\top}]^{l}PB_{w}\right\|_{2\to 2}\mbox{ and }b:=\max_{j}\|w_{j}\|_{2}$
The assumption $w\in\ell_{2}$ implies that $b$ is finite. Moreover, $\hat{A}$ is a Schur matrix and hence $a_{p}\to 0$ as $p\to\infty$. This implies that $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\to 0$ uniformly in $t$ as $p\to\infty$.
@@
The convergence $J_{\infty,p}^{*}\to J_{\infty,nc}^{*}$ follows because $\hat{A}$ is a Schur matrix. Finally, define $X:=\sum_{j=0}^{\infty}(\hat{A}^{\top})^{j}PB_{w}B_{w}^{\top}P\hat{A}^{j}$. $X$ is the solution of the discrete-time Lyapunov equation (58) (See Section 21.1 in [30]). This yields the expression for the optimal cost in (57). ∎
@@ VI Example
Figure 2: Comparison of IH cost versus preview horizon $p$.
Figure 2 shows the optimal IH cost versus the preview horizon $p$ (blue solid). The cost achieved by the finite preview controller is computed using the expression in Equation 47. The plot also shows the cost achieved by the Linear Quadratic Regulator (LQR) state feedback: $u_{t}=-K_{x}x_{t}$ where $K_{x}=H^{-1}B_{u}^{\top}PA$ is the same gain as used in the finite preview controller. The LQR state feedback achieves a cost equal to trace$[PB_{w}B_{w}^{\top}]\approx 33.2$ (red circle). The full information controller $u_{t}=-K_{x}x_{t}-K_{w}w_{t}$ corresponds to the case of $p=0$. This improves on the LQR performance and reduces the cost to $\approx 29.4$. Adding disturbance preview $(p>1)$ further decreases the cost. In fact, the cost monotonically decreases with increasing preview.
@@
Figure 3: Comparison of IH relative error on a log scale versus preview horizon $p$.

### v4

@@ I Introduction
[…]
@@ II Notation
Let $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times m}$ denote the sets of real $n\times 1$ vectors and $n\times m$ matrices, respectively. The superscript $\top$ denote the matrix transpose. Moreover, if $M\in\mathbb{R}^{n\times n}$ then $M\succ 0$ denotes that $M$ is a symmetric positive definite matrix. Similarly, we use $\succeq$, $\prec$, and $\preceq$ for positive semidefinite, negative definite, and negative semidefinite matrices, respectively. A matrix $A\in\mathbb{R}^{n\times n}$ is said to be Schur if the spectral radius is <1<1. We also introduce random variables in the problem formulation and use $\mathbb{E}$ to denote the expectation. Finally, we denote the Kronecker delta $\delta_{ij}$ where $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise.
@@ III Problem Formulation
$i_{t}$ $\coloneq\{x_{0},\dots,x_{t},w_{0},\dots,w_{t+p}\}.$ (7)
@@
Consider the LTI plant (6) with AA nonsingular, $(A,B)$ stabilizable and cost function in (60) with $Q\succeq 0$, $R\succ 0$ and $(A,Q)$ detectable. Find a sequence of policies $K\coloneq\{K_{0},K_{1}\dots\}$ to stabilize the plant and solve:
@@ IV Principle of Optimality
$W_{T}(i_{T})$ $\coloneq\mathbb{E}[g_{T}(\boldsymbol{x}_{T})\,|\,i_{T}],$
$W_{t}(i_{t})$ $\coloneq\min_{u}\mathbb{E}\left[g_{t}(x_{t},u_{t})+W_{t+1}(i_{t+1})\,\middle|\,i_{t},u_{t}=u\right]$
 $\hphantom{:}\quad\mbox{ for }t=0,\dots,T-1.$
@@
[…]
@@
The policy that achieves $J^{*}(i_{0})$ from will be deterministic if we use this choice for $t=0$ to T1T-1. ∎
@@ V-A FH Stochastic LQR with Preview
$K_{x,t}$ $\coloneq H_{t}^{-1}B_{u,t}^{\top}P_{t+1}A_{t},$ (22)
$K_{w,t}$ $\coloneq H_{t}^{-1}B_{u,t}^{\top}P_{t+1}B_{w,t},$ (23)
$K_{v,t}$ $\coloneq H_{t}^{-1}B_{u,t}^{\top}.$ (24)
@@ V-B IH Stochastic LQR with Preview
$K_{x}$ $\coloneq H^{-1}B_{u}^{\top}PA,$ (37)
$K_{w}$ $\coloneq H^{-1}B_{u}^{\top}PB_{w},$ (38)
$K_{v}$ $\coloneq H^{-1}B_{u}^{\top}.$ (39)
@@
Next, use the definition of $u_{t}^{*}$ in (40) to replace $K_{x}x_{t}$ with $u_{t}^{*}-K_{w}w_{t}-K_{v}v_{t+1}$. The per-step cost is thus given by
@@
[…]
@@ V-C Relation to Optimal Noncausal Controller
$i_{t}$ $\coloneq\{x_{0},\dots,x_{t},w_{0},w_{1},\ldots\}.$ (48)
@@
Consider the LTI plant (6) with AA nonsingular, $(A,B)$ stabilizable and cost function in (49) with $Q\succeq 0$, $R\succ 0$ and $(A,Q)$ detectable. Also assume $w\in\ell_{2}$. Find a sequence of deterministic policies $K\coloneq\{k_{0},k_{1}\dots\}$ to stabilize the plant and solve:
@@
This result, as stated, is a special case of Theorem 1 in [34]. It was stated in [34] for a more general LQR cost including a cross term $2x_{t}^{\top}Sw_{t}$. [^4] The proof given in [34] was for signals defined for $t$ from -$\infty$ to $\infty$. However, the optimal noncausal controller is the same for signals defined from $t=0$ to $\infty$. Note that the full preview design (Problem 3) was stated with a deterministic formulation while finite preview design (Problem 2) was derived with stochastic formulation. However, we show next that the controller with finite preview $p<\infty$ converges, in the limit as $p\to\infty$, to the optimal noncausal controller for any fixed disturbance $w\in\ell_{2}$.
@@
Next, we show that $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\to 0$ uniformly in $t$ as $p\to\infty$. Let $x_{t}^{nc}$ and $x_{t}^{p}$ denote the response of the LTI system (6) with some initial condition $x_{0}$ and disturbance $w\in\ell_{2}$ using the non-causal and $p$-step finite preview controller, respectively. The difference between the two commands is:
$u_{t}^{nc}-u_{t}^{p}=-K_{x}(x_{t}^{nc}-x_{t}^{p})-K_{v}(v_{t+1}^{nc}-v_{t+1}^{p}).$ (55)
Next, note that the states satisfy the following recursion:
$x_{t+1}^{nc}-x_{t+1}^{p}$ $=A\,(x_{t}^{nc}-x_{t}^{p})+B_{u}\,(u_{t}^{nc}-u_{t}^{p})$
$=\hat{A}(x_{t}^{nc}-x_{t}^{p})-B_{u}K_{v}\,(v_{t+1}^{nc}-v_{t+1}^{p}),$
where we have used (55) to substitute for $u_{t}^{nc}-u_{t}^{p}$. Iterating from $x_{0}^{nc}-x_{0}^{p}=0$ (since the initial conditions match) gives:
$x_{t}^{nc}-x_{t}^{p}$ $=-\sum_{j=1}^{t}\hat{A}^{t-j}B_{u}K_{v}(v_{j}^{nc}-v_{j}^{p}).$
Substitute this back into (55) to express the difference in the control commands as follows:
$\begin{split}u_{t}^{nc}-u_{t}^{p}=&-K_{v}(v_{t+1}^{nc}-v_{t+1}^{p})\\ &+K_{x}\sum_{j=1}^{t}\hat{A}^{t-j}B_{u}K_{v}(v_{j}^{nc}-v_{j}^{p}).\end{split}$ (56)
To bound the right side, we use (53) to rewrite $(v_{t+1}^{nc}-v_{t+1}^{p})$ in terms of the disturbance:
$v_{t+1}^{nc}-v_{t+1}^{p}=\sum_{l=p+1}^{\infty}(\hat{A}^{\top})^{l}PB_{w}w_{l+t}.$ (57)
Thus, $\|v_{t+1}^{nc}-v_{t+1}^{p}\|_{2}\leq a_{p}\cdot b$ for $t\geq 0$, where
$a_{p}:=\left\|\sum_{l=p+1}^{\infty}[\hat{A}^{\top}]^{l}PB_{w}\right\|_{2\to 2}\mbox{ and }b:=\max_{j}\|w_{j}\|_{2}.$
The assumption $w\in\ell_{2}$ implies that $b$ is finite. Moreover, $\hat{A}$ is a Schur matrix and hence $a_{p}\to 0$ as $p\to\infty$. [^6] The sum $\sum_{l=p+1}^{\infty}[\hat{A}^{\top}]^{l}$ converges to $(I-\hat{A}^{\top})^{-1}\,(\hat{A}^{\top})^{p+1}$. Thus, $a_{p}\leq\|(I-\hat{A}^{\top})^{-1}\|_{2\to 2}\cdot\|(\hat{A}^{\top})^{p+1}\|_{2\to 2}\cdot\|PB_{w}\|_{2\to 2}.$ We have $(\hat{A}^{\top})^{p+1}\to 0$ as $p\to\infty$ and hence $a_{p}\to 0$.
Finally, use this bound and Equation 56 to bound the difference in control commands:
$\|u_{t}^{nc}-u_{t}^{p}\|_{2}\leq a_{p}b\left[c_{1}+c_{2}\cdot\sum_{j=0}^{t-1}\|\hat{A}^{j}\|_{2\to 2}\right],$ (58)
where $c_{1}:=\|K_{v}\|_{2\to 2}$ and $c_{2}:=\|K_{x}\|_{2\to 2}\cdot\|B_{u}K_{v}\|_{2\to 2}$ are finite constants. We can choose any $\alpha\in(\rho(\hat{A}),\,1)$ and $\exists\,T$ such that $\left\|\hat{A}^{j}\right\|\leq\alpha^{j}$ for all $j\geq T$. This follows from Gelfand’s formula (Corollary 5.6.14 in [horn2013matrix]). Therefore, the sum $\sum_{j=0}^{\infty}\|\hat{A}^{j}\|_{2\to 2}$ converges to some finite number $c_{3}$. Thus we can bound the difference in control commands as:
$\|u_{t}^{nc}-u_{t}^{p}\|_{2}\leq a_{p}b\,(c_{1}+c_{2}\,c_{3}),$ (59)
where $b$, $c_{1}$, $c_{2}$, and $c_{3}$ are all finite constants. Moreover $a_{p}\to 0$ as $p\to\infty$. This implies that $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\to 0$ uniformly in $t$ as $p\to\infty$.
@@
The convergence $J_{\infty,p}^{*}\to J_{\infty,nc}^{*}$ follows because $\hat{A}$ is a Schur matrix. Finally, define $\sum_{j=0}^{\infty}(\hat{A}^{\top})^{j}PB_{w}B_{w}^{\top}P\hat{A}^{j}. XX$ is the solution of the discrete-time Lyapunov equation (62) (See Section 21.1 in [30]). This yields the expression for the optimal cost in (61). ∎
@@ VI Example
Figure 2: Comparison of IH cost versus preview horizon pp.
[…]
@@
Figure 3: Comparison of IH relative error on a log scale versus preview horizon pp.